13. A stock has a beta of 1.2 and an expected return of 17 percent. A T-bill earns a rate of 5.1 percent. The beta of a portfolio comprised of these two assets is 0.85. What percentage of the portfolio is invested in the stoc? (1 point) Hint: T-bill eams risk free rate A. 71% B. 77% C.84% D. 89% E. 92% 14. You own a portfolio equally invested in a risk-free asset and two stocks. One of the stocks has a beta of 1.9 and the total portfolio is equally as risky as the market. What is the beta of the second stock? (1 point) A. 0.75 B. 0.80 C. 0.94 D. 1.00 E. 1.10
Answer #1

Answer to Question 13:

Beta of Stock = 1.2
Beta of T-bills = 0
Beta of Portfolio = 0.85

Let weight of stock be x and weight of T-bill be (1-x)

0.85 = x * 1.2 + (1-x) * 0
0.85 = 1.2 * x
x = 0.71
x = 71%

So, weight of stock is 71%

Answer to Question 14:

Beta of Stock 1 = 1.9
Beta of Stock 2 = x
Beta of T-bills = 0
Beta of Portfolio = 1

Weight of stock 1 = Weight of stock 2 = Weight of risk-free asset = 1/3

1 = (1/3) * 1.9 + (1/3) * x + (1/3) * 0
3 = 1.9 + x
x = 1.10

So, beta of stock 2 is 1.10